A car leaves Columbia at 1:00pm traveling at 45mph. A second car leaves Columbia at 2:00 traveling at 50mph. How long will it take the second car to overtake the first car?
Answer by mananth(12270)

Car leaves Columbia at 1.00 pm 45 mph
second car leaves at 2.00pm 50 mph
..
In 1 hour first car has traveled 45 miles.
..
They are travelling in the same direction so effective speed = 50-45 = 5 mph.
every hour second car makes up 5 miles.
distance is 45 miles
..
time = d/r
t=45/5
t = 9 hours
..
m.ananth@hotmail.ca